OpenStudy (anikate):

14,17,22,2,22,n mean=16

OpenStudy (anonymous):

see u hv six observations out of which one is not given but denoted by n sum of all observations Mean = --------------------------- total number of observations we can say Mean * total number of observations = sum of all observations

OpenStudy (anikate):

what?

OpenStudy (anikate):

i'm just 7th grade

OpenStudy (anonymous):

(14+17+22+2+22+n)/6 = 16 Solve for n.

OpenStudy (anikate):

yup

OpenStudy (anonymous):

here mean = 16 number of observations = 6 so using above formula, sum of observations = 16 * 6 = 96 also 14 + 17 + 22 + 2 + 22 + n = 77 + n 77 + n = 96 n = 96 - 77 n = 19

OpenStudy (anikate):

where did 77 come from?

OpenStudy (anonymous):

by adding the numbers..

OpenStudy (anikate):

o

OpenStudy (anonymous):

hv u understood it now????